Prove: If n and a are positive integers and n=(a^2+ 1)/2, then n is the sum of the squares of two consecutive integers (that is, n=k^2+ (k+1)^2 for some integer k).
Answer #1

2 criven n=a²+1 . we want find two Consecutive integens k and kto such that K+ (+1)= n K²+k+1)² = 9211 - o Now we solve equat